Probability and Statistics II GSIA, Carnegie Mellon University 45-734, Spring 2002

Homework #2, due Thursday, March 28, 2002

Please attach relevant eview output to your assignment!

Please do the following problems from the book:

pg 458 27 pg 468-70 31, 33, 43

Also, please use the wine data to:

• Estimate a model to test the theory that vintage affects the price of wine. Use the data from the website with eviews. Test the hypothesis that vintage does not affect price at the 5% level.

• Estimate a 95% confidence interval for the effect of aging a wine one year on its price.